Comp 210 Information Sheet

Fall 1999, MWF 10:00, PL210

Instructional Staff

Lectures:  Kathi Fisler

Office; phone: Duncan 3093; x3834

Email: kfisler@cs.rice.edu

Office Hours: Tues 13:00–14:00, Wed 11:00–12:00 & by appointment

Lab Coord.: John Greiner

Office; phone: Duncan 3118; x3838

Email: greiner@cs.rice.edu

Office Hours: Mon 13:30–15:00, Fri 13:30–15:00 & by appointment

See http://www.owlnet.rice.edu/~comp210/ for a list of TAs, labbies, and updated office hours.

Texts

Required: How to Design Programs, Felleisen et al.

(available on-line and at the campus bookstore)

Optional: The Little Schemer, Friedman & Felleisen

The Seasoned Schemer, Friedman & Felleisen Also see the Web page

Information Sources

The course has a homepage on the World Wide Web:

                           http://www.owlnet.rice.edu/~comp210/

as well as a newsgroup, rice.owlnews.comp210. The first lab will introduce you to a web-browser so that you can access both these sources of info. The homepage contains a syllabus, assignments, (partial) lecture notes, and pointers to other things. The newsgroup is a forum for discussion and also where homework clarifications can be posted. You should read the newsgroup every day or two.

Goals

Your goal in this course is to learn the fundamental, enduring principles of computing and programming. Your ungoal is to get a good grade. A good grade should be the result of learning something; it cannot be your primary motivation in any college course.

Laboratory Sections

There are weekly lab sections held in Ryon 102 (times: T 13:00-14:20, T 14:30-15:50, W 14:00-15:20, W 15:30-16:50, W 19:00-20:20, Th 13:00-14:20, Th 14:30-15:50). Although attendance is not taken, you are responsible for all information discussed in lab. Labs serve two purposes:

• to supplement the lecture material with examples and some additional information, and

• to show how to use various programs and tools to complete the homework.

Therapy Sessions

Weekly (optional) therapy sessions will be held in the evenings at four of the colleges. Therapy sessions provide group study time, with a labby present to answer questions. Graded homeworks will be returned at the therapy sessions.

Computing and Programming Environment

You must have access to the OwlNet Computer Laboratory, a collection of computers and X-terminals located in Ryon 102, Mechanical Engineering 241, Abercrombie A121, Abercrombie B223, and the dorms. Apply for a computer account by typing apply followed by return at any OwlNet computer or terminal. Reply to all the nasty questions and read the instructions at the end of the sessions. You are welcome to use your own computer but we cannot provide technical support if you do. You can download the primary courseware—DrScheme—via a link from the course homepage.

Grades

Your grade for this course will be based on: homework (30%), three exams (10%, 25%, 25%, respectively), and one final project (10%). Makeup exams will only be granted in emergency situations.

Homework

There will be one homework assignment per week. The purpose of the homework is to deepen your understanding of the class and lab material and to prepare you for the exams. Hard-copies of homeworks will be due at the beginning of class (usually on Wednesdays); the homework guidelines on the course web page describe the expected format for submitted assignments. LATE HOMEWORK WILL NOT BE ACCEPTED without prior arrangement. Graded homeworks may be picked up during therapy sessions.

Honor Code Policy

Exams test your individual understanding and knowledge of the material.

Collaboration during exams is not tolerated.

Homeworks help you understand the material and prepare you for the exam. You are encouraged to discuss the homework problems with the lecturers, teaching assistants, lab assistants, and other students, including 210 graduates. You may collaborate on a homework solution with up to one other student; turn in a single joint homework solution with the names of both students on the cover sheet. Your name should appear on exactly one submitted assignment per week: do not submit separate homeworks with different collaborators. All collaborators are responsible for knowing all the submitted material. Warning: The exams are designed to test your understanding of the homework material. If you fail to understand the homework solutions, you cannot understand the lecture material and you cannot pass the course.

Disabilities

Any student with a disability requiring accommodations in this class is encouraged to contact the instructor after class or during office hours. Additionally, students can contact the Coordinator for Disabled Student Services in the RMC Cloisters. If you need accommodations for exams, you must speak with the instructor during the first two weeks of class.